### v1

# User Experience in Dataset Search Platform Interfaces

Abstract.
This research investigates User Experience (UX) issues in dataset search platform interfaces, targeting Google Dataset Search and data.europa.eu. It focuses on 6 areas within UX: Initial Interaction, Search Process, Dataset Exploration, Filtering and Sorting, Dataset Actions, and Assistance and Feedback. The evaluation method combines ’The Pandemic Puzzle’ user task, think-aloud methods, and demographic and post-task questionnaires. 29 strengths and 63 weaknesses were collected from 19 participants involved in roles within technology firm or academia. While certain insights are specific to particular platforms, most are derived from features commonly observed in dataset search platforms across a variety of fields, implying that our findings are broadly applicable. Observations from commonly found features in dataset search platforms across various fields have led to the development of 10 new design prototypes. Unlike literature retrieval, dataset retrieval involves a significant focus on metadata accessibility and quality, each element of which can impact decision-making. To address issues like reading fatigue from metadata presentation, inefficient methods for results searching, filtering, and selection, along with other unresolved user-centric issues on current platforms. These prototypes concentrate on enhancing metadata-related features. They include a redesigned homepage, an improved search bar, better sorting options, an enhanced search result display, a metadata comparison tool, and a navigation guide. Our aim is to improve usability for a wide range of users, including both developers and researchers.
User study, User interaction, Interview, Think-aloud, Search interface, Search platform, Dataset search, Dataset retrieval, Dataset, Metadata, Structured data, Information search, Information retrieval, Google dataset search, data.europa.eu
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; 2024; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Information retrievalccs: Human-centered computing Human computer interaction (HCI)$c$s: Human-centered computing User studiesccs: Information systems Search interfacesccs: Human-centered computing Accessibilityccs: Human-centered computing Empirical studies in HCIccs: Information systems Web searching and information discovery
1. Introduction
In the present data-driven landscape, sectors such as healthcare and social sciences are increasingly reliant on diverse datasets (Benjelloun et al., 2020; Mayer-Schönberger and Cukier, 2013). The sheer volume and complexity of available datasets often result in data silos, hampering their effective usage (Roh et al., 2019; Stonebraker et al., 2018). This dichotomy between the surfeit of publicly available datasets and the particular datasets users can actually locate constitutes a significant challenge (Chapman et al., 2020).
Earlier studies have delved into the intricacies of dataset search systems, highlighting the critical role of metadata and the need for its standardisation (Chapman et al., 2020; Benjelloun et al., 2020). Research has also been conducted on how Google Dataset Search leverages metadata to organise and rank search results (Brickley et al., 2019). Additionally, some focus has been given to understanding user behaviour on governmental dataset platforms, identifying the primary factors that influence search queries (Kacprzak et al., 2019). However, there still exists a void in understanding the user experience in navigating these dataset search platforms, an area this research aims to explore.
For this investigation, Google Dataset Search and data.europa.eu were selected due to their advanced metadata-based search features and high prevalence among researchers. This study assesses user experience in both document retrieval and dataset retrieval, examining existing features of Google Dataset Search and data.europa.eu platforms. Researchers synthesis these elements to pinpoint six aspects of user experience in dataset search. A user task, termed ”The Pandemic Puzzle,” is then designed to explore currently trending areas in dataset search. This study adopts both Concurrent and Stimulated Retrospective think-aloud techniques to investigate user interactions with these dataset retrieval platforms. To enhance the reliability of the think-aloud methods, Thought Prompts and preliminary training sessions are included. Additional context and a swift assessment of perceived usability are provided through demographic details and post-task questionnaires.
This research contributes to enhancing user experience on dataset search platforms in several ways. Firstly, the redesigned homepage emphasizes the platform’s scope, clearly showing content that users are most interested in, aiding them in determining if the site meets their data needs. Secondly, the search bar extends beyond datasets to include non-dataset items like help documents and authors, while providing visual aids for Boolean operators. These operators are effective for leveraging metadata, significantly improve the efficiency of finding relevant datasets. Thirdly, the dataset preview panel using color-blind-friendly colors to represent metadata quality and displaying the most sought-after metadata to reduce reading fatigue and hasten dataset filtering. Fourthly, simplified filters with a display of the remaining dataset count prevent over-filtering. Fifthly, a metadata comparison feature utilizes colors to differentiate quality between two datasets on the same metadata, speeding up the final selection process. Lastly, a visual tutorial clarifies the user interface functions, offering guidance for new users.
These research findings serve as a basis for improving dataset search platforms. The results guide platform developers in understanding user behavior and preferences, aiding in the creation of more efficient dataset discovery tools.
2. Related Works
Information seeking behaviour: in web searches and digital libraries, information seeking behaviour is a well-studied area with established models that map out the process and strategies users follow  (Ellis, 1989). However, the suitability of these models for dataset retrieval is uncertain since they were originally intended for document searches  (Ellis, 1989). Searches are generally of two types: ’lookup’, which is simple and well-supported, and ’exploratory’, which is more complex, involving tasks that require users to learn and discover new information  (Marchionini, 2006). Exploratory searches can be open-ended, leading to uncertainty when information is incomplete  (Marchionini, 2006). The complexity of exploratory searches is especially significant in dataset retrieval, which has become more challenging with initiatives like GO FAIR that encourage the sharing and reuse of datasets. Such initiatives have led to a greater volume and variety of data, highlighting the need for dataset search systems to be designed with the user’s needs in mind, helping them navigate the complex process of finding the right datasets  (Kern and Mathiak, 2015).
Nature of datasets: Chapman et al.  (Chapman et al., 2020) describe a dataset as an ordered collection of either numerical or textual data points. Metadata, crucial for search enhancement and contextualisation, includes elements like data source, format, and collection time. Their work identifies two core dataset search approaches: ’basic’ and ’user-organised.’ The ’basic’ search is found on platforms such as Figshare (Thelwall and Kousha, 2016) and Google Dataset Search (Brickley et al., 2019), where datasets are pre-curated and stored in repositories. The ’user-organised’ search takes place in environments like data lakes (Gao et al., 2018) and data markets (Grubenmann et al., 2018), where users assemble individual data points themselves, as seen in climate studies collecting temperature and humidity metrics in a data lake.
Online Dataset Accessibility: Benjelloun et al. (Benjelloun et al., 2020) examined Schema.org  (sch, 2024), noted for having the largest and most varied corpus of its kind, and found that only 34% of datasets include licensing information, while 44% offer data download links. Additionally, metadata coverage generally falls under 50%, affecting dataset reusability and findability, thereby influencing user experience in dataset searches. Their study also highlights that 37% of datasets lack persistent URLs, underscoring the necessity for stable identifiers like DOIs. Their findings advocate for standardisation in dataset descriptions, licensing, and access. While transitional periods will be required for standardisation, immediate improvements in user experience are achievable.
Mechanism for searching datasets: Brickley et al. (Brickley et al., 2019) describe the dataset search engine’s multi-stage process, starting with web crawlers collecting metadata across the web. This metadata is subsequently standardised and structured into a universal format using frameworks like Schema.org. Cleaned metadata is then matched with a pre-existing database, or a knowledge graph, to add contextual richness to each dataset. The engine assigns unique ’fingerprints’ to datasets to group similar datasets and eliminate redundancies. Brickley et al. also discuss the engine’s use of MapReduce (Brickley et al., 2019) for efficient data processing. Daily scans are performed to capture new or updated datasets. During a search, the engine ranks datasets based on their query relevance, continually updating its methods to improve its effectiveness and reach.
User-centric literature retrieval: from literature  (Ghorab et al., 2013; Kelly et al., 2015; Croft et al., 2010), they summary 10 user considerations: ’Relevance of Results’ focuses on query interpretation accuracy, context sensitivity, and result ordering; ’Ease of Use’ highlights intuitive interfaces and straightforward navigation; ’Performance and Speed’ aims for quick responses and system reliability; ’Search Features’ includes advanced search options and auto-suggestions; ’Quality of Information’ stresses source credibility and data precision; ’Security and Privacy’ emphasizes user consent and encryption; ’Adaptability’ caters to different languages, regions, and devices; ’Presentation of Results’ targets clear labelling and data visualisation options; and ’Personalisation’ utilises user history and preferences for customisation. In addition, ’Accessibility features’, such as voice search and readability, are integrated to ensure inclusive design.
User-centric dataset retrieval: from literature  (Chapman et al., 2020; Kacprzak et al., 2019; Benjelloun et al., 2020), they indicates that factors like search result accuracy, user interface clarity, search speed, data protection, and system adaptability are crucial . Four key areas for tool enhancement emerge: ’Dataset Metadata’ should have comprehensive descriptions, standardisation, frequent updates, and be optimised for search; ’Functionality, Support, & Security’ implies powerful search features, reliable support, robust data security, and wide accessibility; ’Usability’ entails a user-friendly interface, fast performance, and effective visualisation; ’Support’ is augmented by detailed documentation, educational resources, active user communities, and ongoing tool improvements.
Comparing dataset retrieval with traditional literature retrieval: a study  (Kern and Mathiak, 2015) with 53 participants highlighted the necessity for dataset retrieval systems to be tailored to the specific needs of researchers, as their decision-making process is more complex than that for literature retrieval. The study’s key findings emphasized the significance of metadata quality and accessibility, with users showing a willingness to contribute to metadata enhancement. Additionally, the research found that users value supplementary literature for added context and prefer search systems that are intuitive, like those with autocomplete features, and that offer research-relevant updates and suggestions.
User study on dataset search platform: Kacprzak et al. (Kacprzak et al., 2019) scrutinise user interactions on government dataset search platforms, employing both qualitative and quantitative methods. They pinpoint two significant factors influencing user queries: data features and structural layout. Their research indicates user queries often include geospatial and temporal elements. Metrics reveal that search activity peaks during typical working hours, likely indicating professional use. Their study’s data also show most users access platforms via search engines and predominantly use browsers like Chrome, although Internet Explorer retains significant usage, possibly due to its corporate prevalence. They observe that user engagement metrics, such as page views and time spent, increase with return visits, and they identify challenges users face in data discovery as a prospective area for enhancing search functionalities.
3. Approach
The entire section 3 unfolds following the sequential order of the project’s design process. To establish targeted user tasks, researchers first examined all features of the two websites. They then integrated details from user-centered literature and dataset retrieval in the related works to design the final user tasks. Subsequently, corresponding user study methods and recruitment strategies were set up, during which solutions to challenges were proposed.
3.1. Explore Platform Features
Researcher reviews 2 dataset search platforms to document their features and evaluate their functionalities, aiming to create a task for assessing user experience.
3.1.1. Google Dataset Search
Offers 38 core features intended to streamline user interaction and assist data acquisition. Features such as filters for file type, update time, usage rights, and subject categories enable users to personalise search results. Additional functionalities like dataset saving and sharing foster collaborative research. Dataset citation counts are also displayed as a relevance metric. Detailed file information and dataset previews facilitate quick initial evaluations. Query input is enhanced by autocomplete suggestions. Metadata attributes, including identifiers and author details, are visibly displayed. Figure 1 illustrates 15 primary features on the main page.
Figure 1. Google Dataset Search Platform Features
Highlight each specific element of the Google Dataset Search interface with a red box.
Figure 1. Google Dataset Search Platform Features
3.1.2. data.europa.eu
Data.europa.eu expands on Google Dataset Search by offering advanced filtering options, diverse search types, and custom sorting. It allows users to view recent updates, multiple download options, and comprehensive metadata. Quality indicators and similar dataset suggestions are available. User navigation is monitored, and a SPARQL search enables complex queries. Users can view and visualise datasets directly without downloading. Figure 2 illustrates 17 primary features on the main page.
Figure 2. data.europa.eu Platform Features
Highlight each specific element of the data.europa.eu interface with a red box.
Figure 2. data.europa.eu Platform Features
3.2. Explore User Experience Aspects
Following an analysis of both user requirements and platform functionalities, this study delineates six facets of User Experience: Initial Interaction scrutinises the landing page elements like search buttons and language settings for ease of navigation. Search Process evaluates the efficiency of the search mechanisms, considering factors like speed, accuracy, and result relevance. Dataset Exploration focuses on the user-friendliness of accessing individual dataset details, including readability and source information. Usage of Filters and Sorting assesses the relevance and ease of use for sorting and filtering options. Dataset Actions looks at how conveniently users can perform actions like saving and sharing datasets. Finally, Feedback and Help analyses the platform’s support features, emphasising user ease in providing feedback and seeking help.
3.3. Design User Task
This task in Table  1 coined ’The Pandemic Puzzle’ serves as a mechanism to understand user interaction in identifying, assessing, and utilising datasets on COVID-19 and social sciences.
Table 1. The Pandemic Puzzle User Task
User Task Instructions and Considerations
Background
As a health sciences researcher, investigate COVID-19’s impacts on the UK’s life and social sciences sectors. Focus on datasets that are current, freely accessible, commercial-use, well-documented, from reputable sources, and specific to UK regions in tabular form.
1. Initial Interaction
Engage with the platform to locate datasets on ’COVID-19 impact on UK healthcare systems’ and ’social science research during pandemics’. Assess interface usability including visibility of search, delete, and return functions.
2. Search Process
Input specific queries like ’COVID-19 mental health data UK’ and observe platform’s auto-complete suggestions, result speed, and relevance to the UK’s life and social sciences.
3. Dataset Exploration
Scrutinize datasets for relevance, examining names, sources, types, update frequency, geographic specificity, and alignment with the research on UK’s response to COVID-19.
4. Filters and Sorting
Utilize filters to narrow results to recent, open-access datasets, applying sorting to prioritize by relevance or recency. Assess the precision of filtering and sorting mechanisms.
5. Dataset Actions
Identify suitable datasets for longitudinal analysis of the pandemic’s effects, noting options for saving and sharing. Evaluate the efficiency and intuitiveness of these processes.
6. Feedback and Help
Document any usability issues encountered and identify how the platform provides assistance. Explore the process for submitting feedback and seeking help within the platform.
3.4. Use of User Study Methods
3.4.1. Use of Concurrent Think-Aloud
This study uses moderate-complexity tasks, guided by (Kelly et al., 2009). Participants perform tasks that mix quick reactions with thoughtful analysis, such as choosing a relevant dataset and analysing its metadata. These tasks are more involved than basic queries but not as complex as detailed data analysis. This research employs a think-aloud technique as suggested by (Charters, 2003), giving real-time insight into participants’ thinking. VooV software captures both audio and visual data for comprehensive documentation.
3.4.2. Challenges in Concurrent Think-Aloud
One issue is that participants may feel awkward vocalising their thoughts, which can affect natural decision-making and skew results. Participants might also forget to verbalise their thoughts, especially when focused or performing difficult tasks. Another challenge is the risk of cognitive overload due to multitasking, especially in tasks requiring intense concentration (Charters, 2003).
3.4.3. Addressing Challenges During Concurrent Think-Aloud
Pre-Study Training
A practice task requires participants to find a credible COVID-19 vaccination article. This task has four steps: defining the search goal, conducting the search while verbalising, evaluating sources aloud, and choosing an article while explaining the reasons. A reflection stage follows where participants can identify what worked and what needs improvement.
Providing Thought Prompts
Used to help participants articulate their thoughts during this study. These are specific questions or statements aimed at prompting real-time reactions as participants use dataset search platforms. The goal is to get a detailed understanding of the user experience. Examples of thought prompts include: ”What do you think of this dataset after selecting it?” and ”Which filters are you using and why?”.
3.4.4. Use of Stimulated Retrospective Think-Aloud
This research incorporates both Stimulated Retrospective Think-Aloud (SRTA) and Concurrent Think-Aloud (CTA) techniques to achieve a comprehensive view of user interactions (SELA et al., 2000; Charters, 2003). The SRTA method allows participants to reflect on their actions and thoughts, using visual aids or playback footage for reference. Structured SRTA questions guide this reflective process, creating a reliable setting for participants to share their experiences, thereby improving data quality. An example from this study illustrates this approach.
  • Immediate Thought Prompt: ”What are your first impressions of the search results? Are they what you expected?”
  • Possible Immediate Answer: ”There are more results than I anticipated, making it overwhelming. I expected fewer or more targeted outcomes.”
  • Retrospective Thought Prompt: ”Reflecting on the search, do you remember your initial reactions? Did the results meet your expectations?”
  • Possible Retrospective Answer: ”Initially, I was shocked by the sheer volume of information. I now realise my search terms were too broad, leading to the overload. If I were to do it again, I’d use more specific search criteria to narrow down the results.”
In the immediate response, the participant notes the unexpected volume of results. The retrospective answer goes further by identifying the reason for the initial surprise and suggesting a way to improve future searches.
3.5. Implement Questionnaires
3.5.1. Demographic Questionnaires
Utilised at commencement to gain insights into participant backgrounds (Boynton and Greenhalgh, 2004). These questionnaires capture specifics such as professional skills, educational history, and prior familiarity with dataset search utilities. Comprising seven questions, they targets broad categories like ’General’, ’Education’, and ’Occupation’. It solicits information about age group, gender, highest educational attainment, work or study field, frequency of dataset search tool usage, skill level, and previously used tools. Both multiple-choice and open-ended question types are employed.
3.5.2. Post-Task Questionnaires
Issued upon task completion to collect user opinions (Lewis, 2014). They examine usability facets including user satisfaction, perceived utility, and suggested refinements. These surveys encompass four main areas: ’Initial Interaction’, ’Search Operation’, ’Dataset Study’, and ’Filter and Sort Utilisation’. Questions cover a range of topics, from initial impressions and the relevance of search outcomes to the quality of dataset metadata and the efficacy of additional features. A 1-5 rating scale is used for certain queries, while open-ended questions solicit suggestions for enhancements.
3.6. Participant Recruitment
This study recruited 19 participants from a diverse pool, including both academia and industry, via various channels such as LinkedIn, Boss Zhiping, and university email campaigns. This group included 2 UX designers, 2 product managers, 1 special education aide, and 14 students from fields like Human-Computer Interaction and health data science, with ages ranging from 18 to 40. Prior to this study, ethical approval was obtained and informed consent was collected from all participants. All 19 participants tested both tools with the same task.
This study utilized scenarios involving COVID-19 and social sciences in user tasks, primarily to motivate interaction with the platform, rather than to concentrate on the social or health sciences. The goal of this project was to assess overall user experience, focusing on behavior, functionality, and interface design across diverse search websites. These scenarios were instrumental in prompting users to test specific platform features, such as the autocomplete response to ’COVID-19’ in the search bar or the filter effectiveness in the ’social sciences’ category. Consequently, this research was not centered on deriving conclusions specific to social sciences, leading to the recruitment of participants from a wide array of fields.
3.7. Transcription and Coding Scheme
A single researcher manually transcribed 19 videos, organizing the data into 6 primary themes, including ’Initial Interaction’ and ’Search Process’. The coding process combined top-down and bottom-up approaches as outlined by Gu et al. (Gu, 2014). Preconceived notions about web search biases, such as prior beliefs  (White, 2014), informed some of the top-down codes, whereas bottom-up codes emerged and were adapted based on the analysis of participants’ think-aloud sessions, particularly their interactions with interface elements and their engagement with statistics. Codes were designed to be overlapping to capture the multifaceted nature of user experiences. To confirm reliability, the initial coding was reviewed at a later stage by the same researcher, and any deviations were corrected through iterative revisions. Frequency counts were kept for each code to weigh their impact on participant choices.
4. Results
This section is divided into 3 subsections. The first 2 subsections document feedback for Google Dataset Search and data.europe.eu, labeled ’P’ for positive and ’N’ for negative. Each of these subsections includes six sub-subsections corresponding to the user experience elements identified in the user task. All percentage values have been rounded to two decimal places.
While certain insights, such as interface design elements, are specific to each platform, most of the findings regarding mechanisms and functionalities can be applied more broadly to dataset search platforms. Therefore, the last subsection  4.3 presents the development of 10 new design prototypes, based on observations of features common in dataset search platforms across various fields. In the  4.3.3 Dataset Exploration,  4.3.4 Usage of Filters and Sorting,  4.3.5 Dataset Actions, there’s an enhancement in metadata accessibility. For others, the overall current state of dataset search platforms has been improved.
4.1. Google Dataset Search
4.1.1. Initial Interaction
P1: 68% participants noted the search box, search button, help, and feedback features as easily accessible and user-friendly, commenting that the design was ”extremely neat and comfortable.” P2: 79% participants expressed a clear understanding and appreciation of the layout of the search results page. They found the meaning of each filter to be easily comprehensible and praised the filters’ placement at the top of the page, as well as the well-balanced proportions between the preview and detailed information sections. However, N1: 26% participants observed that the language switch button was not easily noticeable, being located at the bottom left corner. One participant said, ”When I first see this Google Dataset Search page, I think of the traditional Google browser search page, where all function buttons are placed at the top of the page and are immediately visible to users.” N2: 16% participants hovered their cursor over the language switch button, indicating confusion. One participant noted, ”Usually such icons are represented by the flag of the country or are explicitly labelled with the language type, which makes it easy to understand at first glance.” N3: In a multitasking scenario involving a note-taking app, 21% participants found that the detailed dataset information appeared compressed and elongated, requiring constant scrolling. They commented, ”The detailed introduction of the dataset is compressed to the point where only two or three words are displayed in one line. This is unreadable.” N4: 16% participants hovered their mouse over the bookmark icon due to its unclear design. One of them stated, ”It looks like a report button.” N5: 11% participants were unsure about making immediate actions like bookmarking, sharing, or citing a dataset. One said, ”I was looking for a way to mark this dataset, but I couldn’t find it,” after extensive scrolling. N6: An undergraduate participant from mainland China pointed out that switching the website’s language only updated the structure, not the content of dataset details. Both the participant and the researchers acknowledged the limitation of the website’s language features for users of browsers lacking built-in translation.
4.1.2. Search Process
P7: 63% participants start to input search queries sequentially, utilising system suggestions before finalising keywords. This strategy aids in objective-specific result retrieval and avoids ineffective keywords. In some cases, system-provided datasets already met research task requirements. A participant noted, ”This feature sometimes yields unexpected results!” P8: Google Dataset Search recognised misspelled words. For instance, typing ”deth” retrieved datasets related to ”death.” P9: For misspelled terms like ”covai19,” Google Dataset Search failed to retrieve datasets but offered suggestions like ”Make sure that all words are spelled correctly.” P10: 21% Participants effectively utilised Boolean and Exact Phrase operators. Use of quotes made a discernable difference in search results. However, N14: Using ”covid-19 cases” as a prefix didn’t yield UK-specific datasets. Switching to ”covid-19 infections” did, but also increased time spent on it. N15: The researchers found that participants began by entering ”England covid-19 cases”, resulting in two system suggestions, with the second one fitting research task requirements. However, not all users were as fortunate. When ”covid-19 cases England” was entered, Google Dataset Search suggested only one dataset, which wasn’t the best match. The rearrangement of search terms had a substantial effect on the results and the time spent searching. N16: 37% participants spent a significant amount of time downloading a dataset to look at its colmun and correlation values, and then gave up on using the dataset because of the large number of missing values in the file. One participant said, ”Luckily the zip file is not too big. N17: Google Dataset Search misunderstood ”covidd-19 infraction” to mean ”covid-19 infection,” revealing its limitations in contextual understanding. The participant commented, ”Google Dataset Search can recognise my spelling errors well, but sometimes it misunderstands what I am trying to convey.” N18: Different spellings of ”covid-19” yielded varying results. The spelling ”covid-19” was more task-related compared to other variants. N19: Misspellings in multiple words made it challenging for users to correct their queries, especially when they didn’t realise a spelling error. N20: Sentence-like queries yielded fewer results than more structured queries. Both aimed to retrieve similar information. N21: The Exclude Words operator ”-” didn’t function as expected, leading to undesired results. N22: Searching for operator-related help was futile. 21% participant remarked, ”The User Guide page of the website is 404, and now it doesn’t provide me with the corresponding help, which is very confusing and wasting my time.”
4.1.3. Dataset Exploration
P13-P16 Participants first pay attention to the title, then the update time and dataset type. One remarked, ”I trust dataset content over the publisher”. Furthermore, dataset ranking influences choices, especially the top three listings. P17 Participants paid particular attention to the descriptions, citation counts, and providers on the right-hand dataset display section. P18 74% participants were indifferent to the specifics of the licence, but 68% preferred its existence. However, N29 53% participants mentioned that they initially cared a lot about the ranking on the left side of the page. 26% participants said they didn’t care about the ranking at all and often checked many datasets. One participant even scrolled to the bottom of the page, causing the page to refresh twice before they stopped checking datasets. One participant stated, ”I feel that both datasets ranked high and low can meet my search requirements”. N30 21% participants were frustrated by incomplete dataset titles and limited preview information. N31 and N32 26% participants showed a lack of interest in the dataset provider and dense description, commenting ”I only care if the dataset has a provider, not the specific provider name” and ”Descriptions too dense to be read only by machines, not people!” N33 37% participants point out a lack of detailed file information, and N34 16% participants indicated a high implementation difficulty in packing more information into the left preview module. One participant said, ”Although I hope to fit more information, I would still prefer to see a clean preview section on a page.” N35 21% participants after exploring a qualifying dataset, would click on the website link given in ’explore at’ to view more information about the dataset. However, 11% participants stated, ”Currently, I have to click on ’Explore at’ to jump to the corresponding page to see more detailed information, which is somewhat troublesome. I need to switch between several websites, and sometimes there are too many tabs, and I forget which page was transferred from which link.”
4.1.4. Usage of Filters and Sorting
P23 68% participants intuitively understood the names and effects of the filters without requiring additional help. N42 32% participants underlined the difficulty in sorting the datasets by their recency; a participant stated, ”The dataset search results are sorted too simply. I just want to find a recent dataset, but there is no way to do it.” N43 26% participants noticed that for multiple choice filters such as ”topic”, when they chose two options, the ”+1” displayed on the filter could be misleading. 16% of them interpret this ”+1” as indicating that only one ”topic” has been selected. N44 16% participants preferred interactive animations for dropdown symbols for a more intuitive interface experience, articulated by a participant’s statement, ”Since this is a dropdown symbol, why doesn’t it have the animation effect of a dropdown symbol? I thought the webpage was not loaded.” N45 16% participants highlighted concerns about the broad time spans in the ’Last updated’ filter; as one participant noted, ”There are few datasets in the last month, and there are too many datasets in the last year. Isn’t there an intermediate time period to choose?” N46 47% participants exposed the inconvenience of filter resets upon search query modification; one participant stated, ”This makes it uncomfortable to use, I have to re-select the filter every time.” N47 21% participants observed that some datasets lacked usage permission tags, causing issues with filter-based searches; a participant asked, ”Now whether I click on ’commercial use allowed’ or ’Non-commercial use allowed’, only two datasets are always displayed, but in fact, there are eight available datasets in the current search results, so what usage permission category do the other six datasets belong to?” N48 11% participants criticised the broad categorisation within the ’Topics’ filter, highlighted by a participant who said, ”Especially the ’Natural Science’ and ’Engineering’ tags, their scope can basically include the scope of other search tags.” N49 47% participants expressed confusion about the classification of the datasets retrieved after applying the filters. One participant stated, ”I don’t even know which filters this dataset satisfies. I need to read the entire introduction to get a feel for it. In this case, the filters don’t seem to have worked and haven’t helped me save time.” Lastly, N50 11% participants ignored filters altogether; one stated, ”Honestly, I ignored the existence of the filters at the beginning, mainly because the first search page of Google Dataset Search doesn’t have filter options, just a search box.”
4.1.5. Dataset Actions
P25 All participants stated that they easily found and used the ”Save”, ”Share”, ”Cite” buttons when they found a suitable dataset. They also reported that it was very simple to locate previously saved datasets through the ”Saved datasets” button. Many participants expressed, ”The position of these three buttons is very clear, and I could see them at first glance”. Yet, N56 16% participants preferred to use their browser’s built-in share function. One participant said, “Compared to directly copying the link from the top of the browser, clicking the share button and then choosing the sharing method seems very complicated.” N57 The researcher asked all participants ”Would you use these four sharing methods? Why?” 37% participants stated they would not use the Facebook and Twitter sharing options. One user stated, “I generally only use the dataset search website in a work context”. Moreover, 21% participants stated, ”I would share this dataset via email, but typically, the dataset is only a small part of our email content. Therefore, I wouldn’t write an email solely because of this dataset; I would reference this dataset where needed in the email, so I wouldn’t use the current email sharing feature.”. N58 21% participants misunderstood the ”Click to copy link” as merely informational text rather than an interactive button, attempting to select the URL beneath it instead. N59 16% participants found that the website does not specify the current citation format and standard. One participant mentioned, “I basically would not directly use the format provided by this citation, because it doesn’t even tell me what format it uses.”
4.1.6. Feedback and Help
P27: 32% participants pointed out that google provides what is a dataset search and how to provide a dataset, which is very friendly to people without basic knowledge. P28: 84% participants easily found the ”Help” and ”Feedback” buttons, and two participants independently browsed the content in the ”Help” page at the beginning. N62: 11% participants found that the information on the user guide page currently shows a ’404 Page not found’ error, and participants need to jump to multiple pages in the help documentation to understand Google Dataset Search, which adds a lot of reading time.
4.2. data.europa.eu
4.2.1. Initial Interaction
P3: 11% participants commended the comprehensive summary of datasets on the initial page of data.europa.eu. One participant cited the inclusion of datasets from 36 countries as instilling confidence, while another was impressed by the sheer number of datasets, stating, ”over 1,000,000 datasets shocked me a lot, I feel this website must contain a suitable dataset for my requirement.” P4: The display of a page path was seen as useful by 26% participants for quick navigation. P5: A window hover and shadow animation were appreciated by 42% participants. One stated, ”compared to Google Dataset Search, this animation makes me feel more clear and conspicuous about what I have chosen.” P6: The visibility of selected filters at the top of the results preview section was praised by 26% participants for providing an overview. One remarked, ”Especially with multi-select filters, I can instantly see all the options I’ve selected, which indeed makes subsequent modifications more convenient for me.” However, N7: 32% Participants using smaller laptop screens criticized the absence of a readily visible search box. One participant stated, ”If I didn’t know at first that this web is a dataset search tool, I wouldn’t believe it has a search bar, because I can’t see it at first glance when I enter this website, and this search engine design style is not common.” N8: 53% participants reported distractions due to unrelated windows. One mentioned, ”I came to search for datasets, not to see any advertisements or events”, and another criticized the switching advertisements, stating, ”The switching always disturbs my thought process when I am thinking about what keywords to put into this search bar.” N9: A UX designer noted the absence of indicators for the search bar, making it appear non-functional. N10: 37% participants reported cumbersome scrolling to select filters, which also shifted positions upon selection. One participant expressed frustration, stating, ”The process of selecting filters is really uncomfortable. I need to scroll the page back and forth many times to complete the selection, and why does the position of the selected filter change?” N11: 11% participants found it inconvenient that text in the preview section couldn’t be copied for note-taking. One said, ”I have no choice but to open this dataset, copy the information I need from its detailed introduction page, which is a waste of time.” N12: Although colour differentiation for tags was appreciated, it was found to be inaccessible for individuals with different types of colour blindness. N13: 11% participants criticized the ’show more’ button as redundant when there’s minimal hidden content. They commented, ”In order to see more content, I need to click the ’show more’ button, but the hidden content is minimal, so I feel this click interaction is meaningless.”
4.2.2. Search Process
P11: 63% participants found the ability to sort datasets by ”Last modified” and ”Last created” beneficial when they had time constraints. One participant stated, ”The feature to sort by release time is very useful, especially after all keywords and filters have been selected, I just want to see the latest published datasets under all the conditions I have set.” P12: 21% participants found the separate search options for catalogues and datasets provided in the search bar to be very effective. One participant stated, ”Catalogues provide me with information about publishers from a specific country, and I can find the dataset I need by searching for keywords in all the datasets published by this publisher, which greatly improves my search efficiency.” N23: 26% participants observed that, unlike Google dataset search, more strings entered in the search box resulted in a broader search scope instead of narrowing it down. Many participants stated, ”I would prefer the search system to narrow down the search scope based on my search strings, rather than continue to expand it.” N24: 16% participants noted that sorting by a single criterion led to extreme results, especially when many datasets remained. One participant stated, ”I hope the priority in dataset searches is often to find the most relevant first and then find the most recent datasets.” N25: 16% participants noticed the absence of transition animations for the dropdown symbols on the right side of ”Datasets” and ”Sort by”. After clicking ”More” or ”Less filter”, the dropdown symbol changes significantly, indicating the unfolding and folding of results. One participant stated, ”This inconsistency in the overall design can mislead me a lot.”
4.2.3. Dataset Exploration
P20 63% participants indicates strong appreciation for the use of colour tags to denote dataset types, as one participant said, ”This design makes me feel very comfortable, and the types of files are clear at a glance.” P21 42% participants found the similar dataset feature useful when the similarity was high, capturing the sentiment with a comment, ”I didn’t expect that this very similar dataset was published even more recently than the dataset I found.” P22 21% participants discusse the utility of selecting ’Items per Page’, elaborating with a participant’s comment, ”I think displaying ten datasets at a time is very few, and I always have to click next page to refresh new datasets, which I don’t want to wait for.” N36 47% participants suggest the possible improvement of using colour tags for keyword matching, as multiple participants said, ”I would prefer these different colour tags to be about keyword information.” N37 37% participants signify the importance of ’Update Time’ and other dataset metadata like size and missing values ratio, summarised by a participant’s comment, ”as these factors influence whether the dataset warrants further examination.” N38 21% participants highlight the emphasis users put on ’Publisher’ and ’Updated’ in the expansion box, but found ’Contact,’ ’Identifiers,’ etc., less significant. One participant noted, ”I don’t need to understand the contact of a dataset at all, my primary purpose is to check whether the metadata of the dataset is suitable.” N39 42% mentions the confusion over non-English keywords and the difficulty in identifying query-related keywords from the list, summarised by one comment, ”This dataset covers too many keywords, and I just want to see at a glance which keywords from my search strings are included.” N40 16% participants expose user difficulty in understanding the ’Less Similar’ tag, with one participant stating, ”Instead of recording similarity here, why not directly state which keywords are included.” N41 16% participants found that datasets with the ”less similarity” tag were generally far from their target dataset. They found that only datasets with the ”similar” tag might be worth further exploration.
4.2.4. Usage of Filters and Sorting
P24 74% participants stated that the dual provision of filtering and sorting methods on the website was lauded for enhancing efficiency in dataset discovery. However, N51 58% participants exposed a time-consuming issue, where participants found that the page refreshes after each filter selection, especially problematic when applying multiple filters. N52 37% participants thought that refreshing the entire page after each filter selection could mislead users into thinking that the filter options are for single selection only, even though some filter options allow multiple selections, potentially resulting in imprecise search results. One participant, after selecting an option and noticing the page refresh and all expanded options collapse, skipped this filter and stated, ”It would be better if it was multi-select,” even though the filter was indeed multi-select. N53 16% participants spotlighted another issue where three participants were unaware that a filter was for single selection until they noticed that their previous selection was overwritten by the new one; a participant stated, ”I thought it was multi-select, and I wouldn’t have known it was single-select if the website’s upper right side didn’t display all my selected filter options.” N54 11% participants noted that the presence of numbers in the text of some filter options, such as ”Version 2”, can occasionally impact their interpretation of the number of datasets. One participant stated, ”At first glance, I thought there were only two datasets that met this filter requirement,” due to the confusion between the dataset count and the numerical information within the filter option’s text. Lastly, N55 26% participants found it easy to overlook the search box within each filter. One participant stated, ”as the placeholder text in the search box says ’Filter’, leading me to assume that it’s a description of the drop-down menu.”
4.2.5. Dataset Actions
P26 42% participants stated that the citation function was clearly visible and comprehensive, offering four different citation styles, which saved them significant time in creating citations. One participant stated, ”This has saved me a lot of time in creating citations.” N60 11% participants highlighted the absence of a direct LaTeX citation option, stating that manual LaTeX citation creation still constitutes a burdensome process. N61 58% participants stated that when they want to compare similarities between datasets, they need to manually copy and paste dataset information into a separate document, which they find cumbersome. One participant stated, ”It would be great if datasets could be compared directly on the website.”
4.2.6. Feedback and Help
P29: 68% participants indicated that they could easily locate the help documentation, and the information related to the functionality was easily discovered and understood. N63: 47% participants observed that while they can easily find general information like ”Contact Us” and ”Documentation”, they are often not interested in other options such as ”Publications” and ”Events” present in the navigation bar when searching for datasets. Their focus is primarily on dataset-related features.
4.3. New Prototypes for Dataset Search Platform
4.3.1. Initial Interaction
We suggest that first impressions are crucial for dataset search platforms. The search bar should be prominently displayed, and icons should be universally recognisable. A language switch button featuring familiar symbols, such as Google Translate, should be easily accessible. Secondary information, including website summaries and brand content, should either be collapsible or relocated to the bottom of the page. Visual cues such as breadcrumb navigation and subtle animations can further improve user engagement and navigation (Shown in Figure  3). These design elements should be consistently implemented across various screen sizes to ensure a user-friendly experience.
Figure 3. First Page of A Dataset Search Tool
Prototype of First Page of Dataset Search Tool
Figure 3. First Page of A Dataset Search Tool
4.3.2. Search Process
To enhance user interaction and platform efficiency, several improvements are recommended. Immediate auto-suggestions and spelling corrections could be provided through real-time prompts like ”did you mean.” Dataset search platform could also incorporate Natural Language Processing techniques such as entity extraction and synonym recognition for more accurate results. Special attention should be paid to varying user needs; for instance, new users might appreciate a help query option in the search bar, while other users may prioritize datasets based on factors like geographic location (Shown in Figure  4). Additionally, implementing common search operators like ”AND,” ”OR,” and ”NOT” can refine search specificity. To make these operators more accessible, they could be visualized in the search bar interface, accompanied by dropdown menus or clickable icons. An ’add more’ option could also be provided for scalable keyword and operator input. This approach not only informs users about advanced search tools but also enables them to generate highly targeted results from the beginning (Shown in Figure  5).
Figure 4. Search Bar Considering Varied User Needs
Prototype of Search Bar Considering Varied User Needs
Figure 4. Search Bar Considering Varied User Needs
Figure 5. Search Bar Considering Enhanced Precision
Prototype of Search Bar Considering Enhanced Precision
Figure 5. Search Bar Considering Enhanced Precision
4.3.3. Dataset Exploration
Dataset previews need to be brief but informative. They should underscore search terms, date of publication, quality of metadata, view counts, licence type, file dimensions, covered regions, and the nature of the dataset. Additionally, licence information should be streamlined to quickly signify if commercial use is permitted (See Figure  6). The previews should also be succinct while offering depth, including expandable segments for extra details. These segments can feature key terms, descriptions of columns, and ratios of missing values, presented in a bullet-point format (See Figure  7).
Figure 6. User-Focused Dataset Previews Section
Prototype of Dataset Preview Considering Prioritizing User Preferences
Figure 6. User-Focused Dataset Previews Section
Figure 7. Expandable Section For Additional Information
Expandable Section For Additional Information about dataset metadata
Figure 7. Expandable Section For Additional Information
Figure 8. Simplified Result Display Section
Prototype of Optimizing Result Previews for Enhanced User Experience
Figure 8. Simplified Result Display Section
Given the numerous dataset previews, users can face reading fatigue. To counteract this, a colour-coded scheme could be employed to highlight datasets meeting user-defined criteria. For example, using dark green to signify criteria that match the user’s definitions, intermediate blue for criteria that don’t match, and pale orang for criteria that the user hasn’t defined (Shown in Figure  8), while being accessible to colour-blind users, tested in Section  4.3.5.
4.3.4. Usage of Filters and Sorting
To enhance the user experience across platforms, several improvements are suggested. Filters should allow both single and multiple selections, symbolised by circles and boxes respectively. These filters should encompass a range of options including time range, dataset size, and license type. Ideally positioned at the top of the page, filters should not necessitate a full page refresh upon each selection. Instead, a subtle bottom-right corner notification should indicate the number of datasets that fit the selected criteria (Shown in Figure  9).
Figure 9. Improved Filter Design
Prototype of Filters Considering Optimizing Placement and Interactivity
Figure 9. Improved Filter Design
Sorting should employ dual criteria: initial sorting based on relevance, followed by secondary sorting according to publication date. These improvements aim for a more efficient and user-friendly dataset search process, with subsequent figures outlining the proposed changes (Shown in Figure  10).
Figure 10. Dual Criteria Sorting
Prototype of Dual Criteria Sorting
Figure 10. Dual Criteria Sorting
4.3.5. Dataset Actions
Improvements in dataset actions could be multifaceted. First, the citation functionalities across platforms should be expanded to include diverse styles like LaTeX, catering to academic requirements. Second, implementing a feature to allow users to directly compare datasets could be invaluable. The attributes of datasets should be colour-coded for more straightforward comparisons. A dark green fill could signify attributes where another dataset performs better than the currently selected one, a intermediate blue fill for areas where the current dataset excels, and a pale orange fill to highlight unique features not offered by the selected dataset. This visual guide would streamline the decision-making process, rendering it more intuitive and less cognitively taxing for users (Shown in Figure  11).
Figure 11. Datasets Comparison Page
Prototype of Datasets Comparison Page
Figure 11. Datasets Comparison Page
The researcher used the Coblis Colour Blindness Simulator (col, 2023) to test 8 types of colour blindness on the Figure  11 provided . The results show clear variations in colour fills for each type of colour blindness, ensuring accessibility for those with colour vision deficiencies (Shown in Figure  12).
Figure 12. Eight Colour Blindness Tests
Coblis Colour Blindness Simulator
Figure 12. Eight Colour Blindness Tests
4.3.6. Feedback and Help
For enhancements, both platforms could incorporate introductory tutorials to familiarize new users with their respective features. A step-by-step guide, necessitating users to acknowledge each tutorial segment before advancing, can ensure a more comprehensive understanding of the platform’s functionalities. By instituting this level of guidance, platforms can increase user engagement, reduce initial learning curves, and subsequently, elevate the overall user experience (Shown in Figure  13).
Figure 13. Tutorial Illustration
Prototype of Tutorial Illustration
Figure 13. Tutorial Illustration
5. Discussion and Limitations
Our study found that users experience significant reading fatigue due to the way metadata is presented on dataset search websites. There’s a mismatch between the metadata displayed on search result previews and what users want to see, increasing the time taken for initial dataset selection. Detailed dataset pages are text-heavy and lack clear solutions for information presentation, with no quick dataset comparison feature. Current dataset websites also suffer from issues like excessive advertisements, irrelevant information, inadequate user interaction feedback, and animations. These platforms have adopted sorting and search methods from literature retrieval platforms, leading to inefficiency in dataset searches with extensive metadata.
To address these issues, our research proposes using distinct color fills to indicate metadata quality, reducing text descriptions to alleviate reading fatigue. Additionally, the color combinations are chosen to be friendly to color-blind users. We also simplify and intuitively display the most crucial metadata in search result previews, facilitating quick initial filtering for users, and an intuitive dataset comparison feature for efficient decision-making, using color fills to differentiate the quality of similar metadata between two datasets. We also suggest search function improvements, emphasizing the use of Boolean operators for effective metadata filtering, and website displays covering scope and functionality to save time in preliminary filtering. These changes aim to increase the efficiency of finding desired datasets.
For limitations, such as user tasks set within a social science context, leading to search results predominantly related to social sciences. However, given the broad scope and abundance of social science datasets, they can be seen as representative of the current state of dataset metadata. Furthermore, since metadata is generally universal and our focus was on accessibility-related user experience, these aspects were generalized in our conclusions, minimizing their impact. Some of our findings, like the improvements for search bar and tutorials, resemble traditional literature search outcomes, but these were already existing issues with current dataset search platforms, they need to be acknowledged as well. Our primary focus was on exploring the accessibility of dataset metadata, specifically how to present it to improve search efficiency, precision, and reduce reading fatigue, addressing key user experience challenges today.
6. Conclusion and Future Works
This research assesses user interactions in dataset search on two platforms: Google Dataset Search and data.europa.eu. Think-aloud protocols and questionnaires are used as methods. Strengths and weaknesses in user experience are identified, particularly in terms of interface design and search functionality related to metadata accessibility. The developed prototypes focus on displays that highlight keywords and metadata, which are of primary interest to users. We also employ color schemes friendly to color-blind users to represent metadata quality, aiming to reduce reading fatigue and improve search efficiency. These features are especially evident in the detailed dataset pages, search results pages, and dataset comparison pages. This study contributes to the field of human-centered information retrieval. In future research, implementing eye-tracking will be a priority to gain more nuanced insights into user behaviour during dataset searches. The inability to employ GazeRecorder on platforms like Google Dataset Search and data.europa.eu suggests the need for alternative eye-tracking technologies. Physical eye-tracking devices could offer a solution to the limitations encountered with software-based approaches, as they are less likely to be hindered by website security protocols.
Acknowledgements.
The author wishes to acknowledge the contributions of supervisors Professor Elena Simperl and Dr. Albert Meroño Peñuela for their insightful guidance and inspiration. Thanks are also directed to Bohui Zhang for editorial assistance. Gratitude is extended to the varied participants of the user study and the Department of Informatics at King’s College London for providing essential research materials and support in participant recruitment.

### v2

@@
-User Experience in Dataset Search Platform Interfaces
-Abstract.
-This research investigates User Experience (UX) issues in dataset search platform interfaces, targeting Google Dataset Search and data.europa.eu.
+User Experience In Dataset Search
+Abstract
+This research investigates User Experience (UX) issues in dataset search, targeting Google Dataset Search and data.europa.eu.
 It focuses on 6 areas within UX: Initial Interaction, Search Process, Dataset Exploration, Filtering and Sorting, Dataset Actions, and Assistance and Feedback.
@@
 Our aim is to improve usability for a wide range of users, including both developers and researchers.
-User study, User interaction, Interview, Think-aloud, Search interface, Search platform, Dataset search, Dataset retrieval, Dataset, Metadata, Structured data, Information search, Information retrieval, Google dataset search, data.europa.eu
-copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; 2024; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Information retrievalccs: Human-centered computing Human computer interaction (HCI)$c$s: Human-centered computing User studiesccs: Information systems Search interfacesccs: Human-centered computing Accessibilityccs: Human-centered computing Empirical studies in HCIccs: Information systems Web searching and information discovery
-1. Introduction
-In the present data-driven landscape, sectors such as healthcare and social sciences are increasingly reliant on diverse datasets (Benjelloun et al., 2020; Mayer-Schönberger and Cukier, 2013).
-The sheer volume and complexity of available datasets often result in data silos, hampering their effective usage (Roh et al., 2019; Stonebraker et al., 2018).
-This dichotomy between the surfeit of publicly available datasets and the particular datasets users can actually locate constitutes a significant challenge (Chapman et al., 2020).
-Earlier studies have delved into the intricacies of dataset search systems, highlighting the critical role of metadata and the need for its standardisation (Chapman et al., 2020; Benjelloun et al., 2020).
-Research has also been conducted on how Google Dataset Search leverages metadata to organise and rank search results (Brickley et al., 2019).
-Additionally, some focus has been given to understanding user behaviour on governmental dataset platforms, identifying the primary factors that influence search queries (Kacprzak et al., 2019).
+Keywords:
+User study User interaction Interview Think-aloud Search interface Search platform Dataset search Dataset retrieval Dataset Metadata Structured data Information search Information retrieval Google dataset search data.europa.eu
+1 Introduction
+In the present data-driven landscape, sectors such as healthcare and social sciences are increasingly reliant on diverse datasets [3, 20].
+The sheer volume and complexity of available datasets often result in data silos, hampering their effective usage [21, 23].
+This dichotomy between the surfeit of publicly available datasets and the particular datasets users can actually locate constitutes a significant challenge [6].
+Earlier studies have delved into the intricacies of dataset search systems, highlighting the critical role of metadata and the need for its standardisation [6, 3].
+Research has also been conducted on how Google Dataset Search leverages metadata to organise and rank search results [5].
+Additionally, some focus has been given to understanding user behaviour on governmental dataset platforms, identifying the primary factors that influence search queries [14].
 However, there still exists a void in understanding the user experience in navigating these dataset search platforms, an area this research aims to explore.
@@
 Researchers synthesis these elements to pinpoint six aspects of user experience in dataset search.
-A user task, termed ”The Pandemic Puzzle,” is then designed to explore currently trending areas in dataset search.
+A user task, termed "The Pandemic Puzzle," is then designed to explore currently trending areas in dataset search.
 This study adopts both Concurrent and Stimulated Retrospective think-aloud techniques to investigate user interactions with these dataset retrieval platforms.
@@
 The results guide platform developers in understanding user behavior and preferences, aiding in the creation of more efficient dataset discovery tools.
-2. Related Works
-Information seeking behaviour: in web searches and digital libraries, information seeking behaviour is a well-studied area with established models that map out the process and strategies users follow  (Ellis, 1989).
-However, the suitability of these models for dataset retrieval is uncertain since they were originally intended for document searches  (Ellis, 1989).
-Searches are generally of two types: ’lookup’, which is simple and well-supported, and ’exploratory’, which is more complex, involving tasks that require users to learn and discover new information  (Marchionini, 2006).
-Exploratory searches can be open-ended, leading to uncertainty when information is incomplete  (Marchionini, 2006).
+2 Related Works
+Information seeking behaviour: in web searches and digital libraries, information seeking behaviour is a well-studied area with established models that map out the process and strategies users follow  [9].
+However, the suitability of these models for dataset retrieval is uncertain since they were originally intended for document searches  [9].
+Searches are generally of two types: ’lookup’, which is simple and well-supported, and ’exploratory’, which is more complex, involving tasks that require users to learn and discover new information  [19].
+Exploratory searches can be open-ended, leading to uncertainty when information is incomplete  [19].
 The complexity of exploratory searches is especially significant in dataset retrieval, which has become more challenging with initiatives like GO FAIR that encourage the sharing and reuse of datasets.
-Such initiatives have led to a greater volume and variety of data, highlighting the need for dataset search systems to be designed with the user’s needs in mind, helping them navigate the complex process of finding the right datasets  (Kern and Mathiak, 2015).
+Such initiatives have led to a greater volume and variety of data, highlighting the need for dataset search systems to be designed with the user’s needs in mind, helping them navigate the complex process of finding the right datasets  [17].
 Nature of datasets: Chapman et al.
-(Chapman et al., 2020) describe a dataset as an ordered collection of either numerical or textual data points.
+[6] describe a dataset as an ordered collection of either numerical or textual data points.
 Metadata, crucial for search enhancement and contextualisation, includes elements like data source, format, and collection time.
-Their work identifies two core dataset search approaches: ’basic’ and ’user-organised.’ The ’basic’ search is found on platforms such as Figshare (Thelwall and Kousha, 2016) and Google Dataset Search (Brickley et al., 2019), where datasets are pre-curated and stored in repositories.
-The ’user-organised’ search takes place in environments like data lakes (Gao et al., 2018) and data markets (Grubenmann et al., 2018), where users assemble individual data points themselves, as seen in climate studies collecting temperature and humidity metrics in a data lake.
+Their work identifies two core dataset search approaches: ’basic’ and ’user-organised.’ The ’basic’ search is found on platforms such as Figshare [24] and Google Dataset Search [5], where datasets are pre-curated and stored in repositories.
+The ’user-organised’ search takes place in environments like data lakes [10] and data markets [12], where users assemble individual data points themselves, as seen in climate studies collecting temperature and humidity metrics in a data lake.
 Online Dataset Accessibility: Benjelloun et al.
-(Benjelloun et al., 2020) examined Schema.org  (sch, 2024), noted for having the largest and most varied corpus of its kind, and found that only 34% of datasets include licensing information, while 44% offer data download links.
+[3] examined Schema.org  [2], noted for having the largest and most varied corpus of its kind, and found that only 34% of datasets include licensing information, while 44% offer data download links.
 Additionally, metadata coverage generally falls under 50%, affecting dataset reusability and findability, thereby influencing user experience in dataset searches.
@@
 Mechanism for searching datasets: Brickley et al.
-(Brickley et al., 2019) describe the dataset search engine’s multi-stage process, starting with web crawlers collecting metadata across the web.
+[5] describe the dataset search engine’s multi-stage process, starting with web crawlers collecting metadata across the web.
 This metadata is subsequently standardised and structured into a universal format using frameworks like Schema.org.
@@
 Brickley et al.
-also discuss the engine’s use of MapReduce (Brickley et al., 2019) for efficient data processing.
+also discuss the engine’s use of MapReduce [5] for efficient data processing.
 Daily scans are performed to capture new or updated datasets.
 During a search, the engine ranks datasets based on their query relevance, continually updating its methods to improve its effectiveness and reach.
-User-centric literature retrieval: from literature  (Ghorab et al., 2013; Kelly et al., 2015; Croft et al., 2010), they summary 10 user considerations: ’Relevance of Results’ focuses on query interpretation accuracy, context sensitivity, and result ordering; ’Ease of Use’ highlights intuitive interfaces and straightforward navigation; ’Performance and Speed’ aims for quick responses and system reliability; ’Search Features’ includes advanced search options and auto-suggestions; ’Quality of Information’ stresses source credibility and data precision; ’Security and Privacy’ emphasizes user consent and encryption; ’Adaptability’ caters to different languages, regions, and devices; ’Presentation of Results’ targets clear labelling and data visualisation options; and ’Personalisation’ utilises user history and preferences for customisation.
+User-centric literature retrieval: from literature  [11, 15, 8], they summary 10 user considerations: ’Relevance of Results’ focuses on query interpretation accuracy, context sensitivity, and result ordering; ’Ease of Use’ highlights intuitive interfaces and straightforward navigation; ’Performance and Speed’ aims for quick responses and system reliability; ’Search Features’ includes advanced search options and auto-suggestions; ’Quality of Information’ stresses source credibility and data precision; ’Security and Privacy’ emphasizes user consent and encryption; ’Adaptability’ caters to different languages, regions, and devices; ’Presentation of Results’ targets clear labelling and data visualisation options; and ’Personalisation’ utilises user history and preferences for customisation.
 In addition, ’Accessibility features’, such as voice search and readability, are integrated to ensure inclusive design.
-User-centric dataset retrieval: from literature  (Chapman et al., 2020; Kacprzak et al., 2019; Benjelloun et al., 2020), they indicates that factors like search result accuracy, user interface clarity, search speed, data protection, and system adaptability are crucial .
+User-centric dataset retrieval: from literature  [6, 14, 3], they indicates that factors like search result accuracy, user interface clarity, search speed, data protection, and system adaptability are crucial .
 Four key areas for tool enhancement emerge: ’Dataset Metadata’ should have comprehensive descriptions, standardisation, frequent updates, and be optimised for search; ’Functionality, Support, & Security’ implies powerful search features, reliable support, robust data security, and wide accessibility; ’Usability’ entails a user-friendly interface, fast performance, and effective visualisation; ’Support’ is augmented by detailed documentation, educational resources, active user communities, and ongoing tool improvements.
-Comparing dataset retrieval with traditional literature retrieval: a study  (Kern and Mathiak, 2015) with 53 participants highlighted the necessity for dataset retrieval systems to be tailored to the specific needs of researchers, as their decision-making process is more complex than that for literature retrieval.
+Comparing dataset retrieval with traditional literature retrieval: a study  [17] with 53 participants highlighted the necessity for dataset retrieval systems to be tailored to the specific needs of researchers, as their decision-making process is more complex than that for literature retrieval.
 The study’s key findings emphasized the significance of metadata quality and accessibility, with users showing a willingness to contribute to metadata enhancement.
@@
 User study on dataset search platform: Kacprzak et al.
-(Kacprzak et al., 2019) scrutinise user interactions on government dataset search platforms, employing both qualitative and quantitative methods.
+[14] scrutinise user interactions on government dataset search platforms, employing both qualitative and quantitative methods.
 They pinpoint two significant factors influencing user queries: data features and structural layout.
@@
 They observe that user engagement metrics, such as page views and time spent, increase with return visits, and they identify challenges users face in data discovery as a prospective area for enhancing search functionalities.
-3. Approach
+3 Approach
 The entire section 3 unfolds following the sequential order of the project’s design process.
@@
 Subsequently, corresponding user study methods and recruitment strategies were set up, during which solutions to challenges were proposed.
-3.1. Explore Platform Features
+3.1 Explore Platform Features
 Researcher reviews 2 dataset search platforms to document their features and evaluate their functionalities, aiming to create a task for assessing user experience.
-3.1.1. Google Dataset Search
+3.1.1 Google Dataset Search
 Offers 38 core features intended to streamline user interaction and assist data acquisition.
@@
 Figure 1 illustrates 15 primary features on the main page.
-Figure 1. Google Dataset Search Platform Features
-Highlight each specific element of the Google Dataset Search interface with a red box.
-Figure 1. Google Dataset Search Platform Features
-3.1.2. data.europa.eu
+Figure 1: Google Dataset Search Platform Features
+3.1.2 data.europa.eu
 Data.europa.eu expands on Google Dataset Search by offering advanced filtering options, diverse search types, and custom sorting.
@@
 Figure 2 illustrates 17 primary features on the main page.
-Figure 2. data.europa.eu Platform Features
-Highlight each specific element of the data.europa.eu interface with a red box.
-Figure 2. data.europa.eu Platform Features
-3.2. Explore User Experience Aspects
+Figure 2: data.europa.eu Platform Features
+3.2 Explore User Experience Aspects
 Following an analysis of both user requirements and platform functionalities, this study delineates six facets of User Experience: Initial Interaction scrutinises the landing page elements like search buttons and language settings for ease of navigation.
@@
 Finally, Feedback and Help analyses the platform’s support features, emphasising user ease in providing feedback and seeking help.
-3.3. Design User Task
+3.3 Design User Task
 This task in Table  1 coined ’The Pandemic Puzzle’ serves as a mechanism to understand user interaction in identifying, assessing, and utilising datasets on COVID-19 and social sciences.
-Table 1. The Pandemic Puzzle User Task
+Table 1: The Pandemic Puzzle User Task
 User Task Instructions and Considerations
-Background
-As a health sciences researcher, investigate COVID-19’s impacts on the UK’s life and social sciences sectors.
+Background As a health sciences researcher, investigate COVID-19’s impacts on the UK’s life and social sciences sectors.
 Focus on datasets that are current, freely accessible, commercial-use, well-documented, from reputable sources, and specific to UK regions in tabular form.
-1. Initial Interaction
-Engage with the platform to locate datasets on ’COVID-19 impact on UK healthcare systems’ and ’social science research during pandemics’.
+1.
+Initial Interaction Engage with the platform to locate datasets on ’COVID-19 impact on UK healthcare systems’ and ’social science research during pandemics’.
 Assess interface usability including visibility of search, delete, and return functions.
-2. Search Process
-Input specific queries like ’COVID-19 mental health data UK’ and observe platform’s auto-complete suggestions, result speed, and relevance to the UK’s life and social sciences.
-3. Dataset Exploration
-Scrutinize datasets for relevance, examining names, sources, types, update frequency, geographic specificity, and alignment with the research on UK’s response to COVID-19.
-4. Filters and Sorting
-Utilize filters to narrow results to recent, open-access datasets, applying sorting to prioritize by relevance or recency. Assess the precision of filtering and sorting mechanisms.
-5. Dataset Actions
-Identify suitable datasets for longitudinal analysis of the pandemic’s effects, noting options for saving and sharing. Evaluate the efficiency and intuitiveness of these processes.
-6. Feedback and Help
-Document any usability issues encountered and identify how the platform provides assistance. Explore the process for submitting feedback and seeking help within the platform.
-3.4. Use of User Study Methods
-3.4.1. Use of Concurrent Think-Aloud
-This study uses moderate-complexity tasks, guided by (Kelly et al., 2009).
+2. Search Process Input specific queries like ’COVID-19 mental health data UK’ and observe platform’s auto-complete suggestions, result speed, and relevance to the UK’s life and social sciences.
+3. Dataset Exploration Scrutinize datasets for relevance, examining names, sources, types, update frequency, geographic specificity, and alignment with the research on UK’s response to COVID-19.
+4.
+Filters and Sorting Utilize filters to narrow results to recent, open-access datasets, applying sorting to prioritize by relevance or recency.
+Assess the precision of filtering and sorting mechanisms.
+5. Dataset Actions Identify suitable datasets for longitudinal analysis of the pandemic’s effects, noting options for saving and sharing. Evaluate the efficiency and intuitiveness of these processes.
+6. Feedback and Help Document any usability issues encountered and identify how the platform provides assistance. Explore the process for submitting feedback and seeking help within the platform.
+3.4 Use of User Study Methods
+3.4.1 Use of Concurrent Think-Aloud
+This study uses moderate-complexity tasks, guided by [16].
 Participants perform tasks that mix quick reactions with thoughtful analysis, such as choosing a relevant dataset and analysing its metadata.
 These tasks are more involved than basic queries but not as complex as detailed data analysis.
-This research employs a think-aloud technique as suggested by (Charters, 2003), giving real-time insight into participants’ thinking.
+This research employs a think-aloud technique as suggested by [7], giving real-time insight into participants’ thinking.
 VooV software captures both audio and visual data for comprehensive documentation.
-3.4.2. Challenges in Concurrent Think-Aloud
+3.4.2 Challenges in Concurrent Think-Aloud
 One issue is that participants may feel awkward vocalising their thoughts, which can affect natural decision-making and skew results.
 Participants might also forget to verbalise their thoughts, especially when focused or performing difficult tasks.
-Another challenge is the risk of cognitive overload due to multitasking, especially in tasks requiring intense concentration (Charters, 2003).
-3.4.3. Addressing Challenges During Concurrent Think-Aloud
+Another challenge is the risk of cognitive overload due to multitasking, especially in tasks requiring intense concentration [7].
+3.4.3 Addressing Challenges During Concurrent Think-Aloud
 Pre-Study Training
@@
 The goal is to get a detailed understanding of the user experience.
-Examples of thought prompts include: ”What do you think of this dataset after selecting it?” and ”Which filters are you using and why?”.
-3.4.4. Use of Stimulated Retrospective Think-Aloud
-This research incorporates both Stimulated Retrospective Think-Aloud (SRTA) and Concurrent Think-Aloud (CTA) techniques to achieve a comprehensive view of user interactions (SELA et al., 2000; Charters, 2003).
+Examples of thought prompts include: "What do you think of this dataset after selecting it?" and "Which filters are you using and why?".
+3.4.4 Use of Stimulated Retrospective Think-Aloud
+This research incorporates both Stimulated Retrospective Think-Aloud (SRTA) and Concurrent Think-Aloud (CTA) techniques to achieve a comprehensive view of user interactions [22, 7].
 The SRTA method allows participants to reflect on their actions and thoughts, using visual aids or playback footage for reference.
@@
 An example from this study illustrates this approach.
-  • Immediate Thought Prompt: ”What are your first impressions of the search results? Are they what you expected?”
-  • Possible Immediate Answer: ”There are more results than I anticipated, making it overwhelming. I expected fewer or more targeted outcomes.”
-  • Retrospective Thought Prompt: ”Reflecting on the search, do you remember your initial reactions? Did the results meet your expectations?”
-  • Possible Retrospective Answer: ”Initially, I was shocked by the sheer volume of information.
+  • Retrospective Thought Prompt: "Reflecting on the search, do you remember your initial reactions? Did the results meet your expectations?"
+  • Possible Retrospective Answer: "Initially, I was shocked by the sheer volume of information.
 I now realise my search terms were too broad, leading to the overload.
-If I were to do it again, I’d use more specific search criteria to narrow down the results.”
+If I were to do it again, I’d use more specific search criteria to narrow down the results."
 In the immediate response, the participant notes the unexpected volume of results.
 The retrospective answer goes further by identifying the reason for the initial surprise and suggesting a way to improve future searches.
-3.5. Implement Questionnaires
-3.5.1. Demographic Questionnaires
-Utilised at commencement to gain insights into participant backgrounds (Boynton and Greenhalgh, 2004).
+3.5 Implement Questionnaires
+3.5.1 Demographic Questionnaires
+Utilised at commencement to gain insights into participant backgrounds [4].
 These questionnaires capture specifics such as professional skills, educational history, and prior familiarity with dataset search utilities.
@@
 Both multiple-choice and open-ended question types are employed.
-3.5.2. Post-Task Questionnaires
-Issued upon task completion to collect user opinions (Lewis, 2014).
+3.5.2 Post-Task Questionnaires
+Issued upon task completion to collect user opinions [18].
 They examine usability facets including user satisfaction, perceived utility, and suggested refinements.
@@
 A 1-5 rating scale is used for certain queries, while open-ended questions solicit suggestions for enhancements.
-3.6. Participant Recruitment
-This study recruited 19 participants from a diverse pool, including both academia and industry, via various channels such as LinkedIn, Boss Zhiping, and university email campaigns.
-This group included 2 UX designers, 2 product managers, 1 special education aide, and 14 students from fields like Human-Computer Interaction and health data science, with ages ranging from 18 to 40.
-Prior to this study, ethical approval was obtained and informed consent was collected from all participants.
-All 19 participants tested both tools with the same task.
-This study utilized scenarios involving COVID-19 and social sciences in user tasks, primarily to motivate interaction with the platform, rather than to concentrate on the social or health sciences.
-The goal of this project was to assess overall user experience, focusing on behavior, functionality, and interface design across diverse search websites.
-These scenarios were instrumental in prompting users to test specific platform features, such as the autocomplete response to ’COVID-19’ in the search bar or the filter effectiveness in the ’social sciences’ category.
-Consequently, this research was not centered on deriving conclusions specific to social sciences, leading to the recruitment of participants from a wide array of fields.
-3.7. Transcription and Coding Scheme
+3.6 Participant Recruitment
+This study involved 19 participants from diverse backgrounds, including academia and industry, recruited through LinkedIn, Boss Zhiping, and university email campaigns.
+This group consisted of 2 UX designers, 2 product managers, 1 special education aide, and 14 students specializing in Human-Computer Interaction and health data science, with ages ranging from 18 to 40.
+Ethical approval was obtained, and informed consent was collected from all participants.
+All participants tested both tools using the same task.
+This study used scenarios related to COVID-19 and social sciences to motivate user interaction with the platforms, focusing on overall user experience, behavior, functionality, and interface design across various search websites.
+The scenarios prompted users to test specific platform features, such as autocomplete responses to ’COVID-19’ and filter effectiveness in the ’social sciences’ category.
+The research did not aim to draw conclusions specific to social sciences, which is why participants were recruited from a wide range of fields.
+3.7 Transcription and Coding Scheme
 A single researcher manually transcribed 19 videos, organizing the data into 6 primary themes, including ’Initial Interaction’ and ’Search Process’.
 The coding process combined top-down and bottom-up approaches as outlined by Gu et al.
-(Gu, 2014).
-Preconceived notions about web search biases, such as prior beliefs  (White, 2014), informed some of the top-down codes, whereas bottom-up codes emerged and were adapted based on the analysis of participants’ think-aloud sessions, particularly their interactions with interface elements and their engagement with statistics.
+[13].
+Preconceived notions about web search biases, such as prior beliefs  [25], informed some of the top-down codes, whereas bottom-up codes emerged and were adapted based on the analysis of participants’ think-aloud sessions, particularly their interactions with interface elements and their engagement with statistics.
 Codes were designed to be overlapping to capture the multifaceted nature of user experiences.
@@
 Frequency counts were kept for each code to weigh their impact on participant choices.
-4. Results
+4 Results
 This section is divided into 3 subsections.
 The first 2 subsections document feedback for Google Dataset Search and data.europe.eu, labeled ’P’ for positive and ’N’ for negative.
-Each of these subsections includes six sub-subsections corresponding to the user experience elements identified in the user task.
+Each of these subsections includes six user experience elements identified in the user task.
 All percentage values have been rounded to two decimal places.
@@
 Therefore, the last subsection  4.3 presents the development of 10 new design prototypes, based on observations of features common in dataset search platforms across various fields.
-In the  4.3.3 Dataset Exploration,  4.3.4 Usage of Filters and Sorting,  4.3.5 Dataset Actions, there’s an enhancement in metadata accessibility.
-For others, the overall current state of dataset search platforms has been improved.
-4.1. Google Dataset Search
-4.1.1. Initial Interaction
-P1: 68% participants noted the search box, search button, help, and feedback features as easily accessible and user-friendly, commenting that the design was ”extremely neat and comfortable.” P2: 79% participants expressed a clear understanding and appreciation of the layout of the search results page.
-They found the meaning of each filter to be easily comprehensible and praised the filters’ placement at the top of the page, as well as the well-balanced proportions between the preview and detailed information sections.
-However, N1: 26% participants observed that the language switch button was not easily noticeable, being located at the bottom left corner.
-One participant said, ”When I first see this Google Dataset Search page, I think of the traditional Google browser search page, where all function buttons are placed at the top of the page and are immediately visible to users.” N2: 16% participants hovered their cursor over the language switch button, indicating confusion.
-One participant noted, ”Usually such icons are represented by the flag of the country or are explicitly labelled with the language type, which makes it easy to understand at first glance.” N3: In a multitasking scenario involving a note-taking app, 21% participants found that the detailed dataset information appeared compressed and elongated, requiring constant scrolling.
-They commented, ”The detailed introduction of the dataset is compressed to the point where only two or three words are displayed in one line.
-This is unreadable.” N4: 16% participants hovered their mouse over the bookmark icon due to its unclear design.
-One of them stated, ”It looks like a report button.” N5: 11% participants were unsure about making immediate actions like bookmarking, sharing, or citing a dataset.
-One said, ”I was looking for a way to mark this dataset, but I couldn’t find it,” after extensive scrolling.
-N6: An undergraduate participant from mainland China pointed out that switching the website’s language only updated the structure, not the content of dataset details.
-Both the participant and the researchers acknowledged the limitation of the website’s language features for users of browsers lacking built-in translation.
-4.1.2. Search Process
-P7: 63% participants start to input search queries sequentially, utilising system suggestions before finalising keywords.
-This strategy aids in objective-specific result retrieval and avoids ineffective keywords.
-In some cases, system-provided datasets already met research task requirements.
-A participant noted, ”This feature sometimes yields unexpected results!” P8: Google Dataset Search recognised misspelled words.
-For instance, typing ”deth” retrieved datasets related to ”death.” P9: For misspelled terms like ”covai19,” Google Dataset Search failed to retrieve datasets but offered suggestions like ”Make sure that all words are spelled correctly.” P10: 21% Participants effectively utilised Boolean and Exact Phrase operators.
-Use of quotes made a discernable difference in search results.
-However, N14: Using ”covid-19 cases” as a prefix didn’t yield UK-specific datasets.
-Switching to ”covid-19 infections” did, but also increased time spent on it.
-N15: The researchers found that participants began by entering ”England covid-19 cases”, resulting in two system suggestions, with the second one fitting research task requirements.
-However, not all users were as fortunate.
-When ”covid-19 cases England” was entered, Google Dataset Search suggested only one dataset, which wasn’t the best match.
-The rearrangement of search terms had a substantial effect on the results and the time spent searching.
-N16: 37% participants spent a significant amount of time downloading a dataset to look at its colmun and correlation values, and then gave up on using the dataset because of the large number of missing values in the file.
-One participant said, ”Luckily the zip file is not too big.
-N17: Google Dataset Search misunderstood ”covidd-19 infraction” to mean ”covid-19 infection,” revealing its limitations in contextual understanding.
-The participant commented, ”Google Dataset Search can recognise my spelling errors well, but sometimes it misunderstands what I am trying to convey.” N18: Different spellings of ”covid-19” yielded varying results.
-The spelling ”covid-19” was more task-related compared to other variants.
-N19: Misspellings in multiple words made it challenging for users to correct their queries, especially when they didn’t realise a spelling error.
-N20: Sentence-like queries yielded fewer results than more structured queries.
-Both aimed to retrieve similar information.
-N21: The Exclude Words operator ”-” didn’t function as expected, leading to undesired results.
-N22: Searching for operator-related help was futile.
-21% participant remarked, ”The User Guide page of the website is 404, and now it doesn’t provide me with the corresponding help, which is very confusing and wasting my time.”
-4.1.3. Dataset Exploration
-P13-P16 Participants first pay attention to the title, then the update time and dataset type.
-One remarked, ”I trust dataset content over the publisher”.
-Furthermore, dataset ranking influences choices, especially the top three listings.
-P17 Participants paid particular attention to the descriptions, citation counts, and providers on the right-hand dataset display section.
-P18 74% participants were indifferent to the specifics of the licence, but 68% preferred its existence.
-However, N29 53% participants mentioned that they initially cared a lot about the ranking on the left side of the page.
-26% participants said they didn’t care about the ranking at all and often checked many datasets.
-One participant even scrolled to the bottom of the page, causing the page to refresh twice before they stopped checking datasets.
-One participant stated, ”I feel that both datasets ranked high and low can meet my search requirements”.
-N30 21% participants were frustrated by incomplete dataset titles and limited preview information.
-N31 and N32 26% participants showed a lack of interest in the dataset provider and dense description, commenting ”I only care if the dataset has a provider, not the specific provider name” and ”Descriptions too dense to be read only by machines, not people!” N33 37% participants point out a lack of detailed file information, and N34 16% participants indicated a high implementation difficulty in packing more information into the left preview module.
-One participant said, ”Although I hope to fit more information, I would still prefer to see a clean preview section on a page.” N35 21% participants after exploring a qualifying dataset, would click on the website link given in ’explore at’ to view more information about the dataset.
-However, 11% participants stated, ”Currently, I have to click on ’Explore at’ to jump to the corresponding page to see more detailed information, which is somewhat troublesome.
-I need to switch between several websites, and sometimes there are too many tabs, and I forget which page was transferred from which link.”
-4.1.4. Usage of Filters and Sorting
-P23 68% participants intuitively understood the names and effects of the filters without requiring additional help.
-N42 32% participants underlined the difficulty in sorting the datasets by their recency; a participant stated, ”The dataset search results are sorted too simply.
-I just want to find a recent dataset, but there is no way to do it.” N43 26% participants noticed that for multiple choice filters such as ”topic”, when they chose two options, the ”+1” displayed on the filter could be misleading.
-16% of them interpret this ”+1” as indicating that only one ”topic” has been selected.
-N44 16% participants preferred interactive animations for dropdown symbols for a more intuitive interface experience, articulated by a participant’s statement, ”Since this is a dropdown symbol, why doesn’t it have the animation effect of a dropdown symbol?
-I thought the webpage was not loaded.” N45 16% participants highlighted concerns about the broad time spans in the ’Last updated’ filter; as one participant noted, ”There are few datasets in the last month, and there are too many datasets in the last year.
-Isn’t there an intermediate time period to choose?” N46 47% participants exposed the inconvenience of filter resets upon search query modification; one participant stated, ”This makes it uncomfortable to use, I have to re-select the filter every time.” N47 21% participants observed that some datasets lacked usage permission tags, causing issues with filter-based searches; a participant asked, ”Now whether I click on ’commercial use allowed’ or ’Non-commercial use allowed’, only two datasets are always displayed, but in fact, there are eight available datasets in the current search results, so what usage permission category do the other six datasets belong to?” N48 11% participants criticised the broad categorisation within the ’Topics’ filter, highlighted by a participant who said, ”Especially the ’Natural Science’ and ’Engineering’ tags, their scope can basically include the scope of other search tags.” N49 47% participants expressed confusion about the classification of the datasets retrieved after applying the filters.
-One participant stated, ”I don’t even know which filters this dataset satisfies.
-I need to read the entire introduction to get a feel for it.
-In this case, the filters don’t seem to have worked and haven’t helped me save time.” Lastly, N50 11% participants ignored filters altogether; one stated, ”Honestly, I ignored the existence of the filters at the beginning, mainly because the first search page of Google Dataset Search doesn’t have filter options, just a search box.”
-4.1.5. Dataset Actions
-P25 All participants stated that they easily found and used the ”Save”, ”Share”, ”Cite” buttons when they found a suitable dataset.
-They also reported that it was very simple to locate previously saved datasets through the ”Saved datasets” button.
-Many participants expressed, ”The position of these three buttons is very clear, and I could see them at first glance”.
-Yet, N56 16% participants preferred to use their browser’s built-in share function.
-One participant said, “Compared to directly copying the link from the top of the browser, clicking the share button and then choosing the sharing method seems very complicated.” N57 The researcher asked all participants ”Would you use these four sharing methods?
-Why?” 37% participants stated they would not use the Facebook and Twitter sharing options.
-One user stated, “I generally only use the dataset search website in a work context”.
-Moreover, 21% participants stated, ”I would share this dataset via email, but typically, the dataset is only a small part of our email content.
-Therefore, I wouldn’t write an email solely because of this dataset; I would reference this dataset where needed in the email, so I wouldn’t use the current email sharing feature.”.
-N58 21% participants misunderstood the ”Click to copy link” as merely informational text rather than an interactive button, attempting to select the URL beneath it instead.
-N59 16% participants found that the website does not specify the current citation format and standard.
-One participant mentioned, “I basically would not directly use the format provided by this citation, because it doesn’t even tell me what format it uses.”
-4.1.6. Feedback and Help
-P27: 32% participants pointed out that google provides what is a dataset search and how to provide a dataset, which is very friendly to people without basic knowledge.
-P28: 84% participants easily found the ”Help” and ”Feedback” buttons, and two participants independently browsed the content in the ”Help” page at the beginning.
-N62: 11% participants found that the information on the user guide page currently shows a ’404 Page not found’ error, and participants need to jump to multiple pages in the help documentation to understand Google Dataset Search, which adds a lot of reading time.
-4.2. data.europa.eu
-4.2.1. Initial Interaction
-P3: 11% participants commended the comprehensive summary of datasets on the initial page of data.europa.eu.
-One participant cited the inclusion of datasets from 36 countries as instilling confidence, while another was impressed by the sheer number of datasets, stating, ”over 1,000,000 datasets shocked me a lot, I feel this website must contain a suitable dataset for my requirement.” P4: The display of a page path was seen as useful by 26% participants for quick navigation.
-P5: A window hover and shadow animation were appreciated by 42% participants.
-One stated, ”compared to Google Dataset Search, this animation makes me feel more clear and conspicuous about what I have chosen.” P6: The visibility of selected filters at the top of the results preview section was praised by 26% participants for providing an overview.
-One remarked, ”Especially with multi-select filters, I can instantly see all the options I’ve selected, which indeed makes subsequent modifications more convenient for me.” However, N7: 32% Participants using smaller laptop screens criticized the absence of a readily visible search box.
-One participant stated, ”If I didn’t know at first that this web is a dataset search tool, I wouldn’t believe it has a search bar, because I can’t see it at first glance when I enter this website, and this search engine design style is not common.” N8: 53% participants reported distractions due to unrelated windows.
-One mentioned, ”I came to search for datasets, not to see any advertisements or events”, and another criticized the switching advertisements, stating, ”The switching always disturbs my thought process when I am thinking about what keywords to put into this search bar.” N9: A UX designer noted the absence of indicators for the search bar, making it appear non-functional.
-N10: 37% participants reported cumbersome scrolling to select filters, which also shifted positions upon selection.
-One participant expressed frustration, stating, ”The process of selecting filters is really uncomfortable.
-I need to scroll the page back and forth many times to complete the selection, and why does the position of the selected filter change?” N11: 11% participants found it inconvenient that text in the preview section couldn’t be copied for note-taking.
-One said, ”I have no choice but to open this dataset, copy the information I need from its detailed introduction page, which is a waste of time.” N12: Although colour differentiation for tags was appreciated, it was found to be inaccessible for individuals with different types of colour blindness.
-N13: 11% participants criticized the ’show more’ button as redundant when there’s minimal hidden content.
-They commented, ”In order to see more content, I need to click the ’show more’ button, but the hidden content is minimal, so I feel this click interaction is meaningless.”
-4.2.2. Search Process
-P11: 63% participants found the ability to sort datasets by ”Last modified” and ”Last created” beneficial when they had time constraints.
-One participant stated, ”The feature to sort by release time is very useful, especially after all keywords and filters have been selected, I just want to see the latest published datasets under all the conditions I have set.” P12: 21% participants found the separate search options for catalogues and datasets provided in the search bar to be very effective.
-One participant stated, ”Catalogues provide me with information about publishers from a specific country, and I can find the dataset I need by searching for keywords in all the datasets published by this publisher, which greatly improves my search efficiency.” N23: 26% participants observed that, unlike Google dataset search, more strings entered in the search box resulted in a broader search scope instead of narrowing it down.
-Many participants stated, ”I would prefer the search system to narrow down the search scope based on my search strings, rather than continue to expand it.” N24: 16% participants noted that sorting by a single criterion led to extreme results, especially when many datasets remained.
-One participant stated, ”I hope the priority in dataset searches is often to find the most relevant first and then find the most recent datasets.” N25: 16% participants noticed the absence of transition animations for the dropdown symbols on the right side of ”Datasets” and ”Sort by”.
-After clicking ”More” or ”Less filter”, the dropdown symbol changes significantly, indicating the unfolding and folding of results.
-One participant stated, ”This inconsistency in the overall design can mislead me a lot.”
-4.2.3. Dataset Exploration
-P20 63% participants indicates strong appreciation for the use of colour tags to denote dataset types, as one participant said, ”This design makes me feel very comfortable, and the types of files are clear at a glance.” P21 42% participants found the similar dataset feature useful when the similarity was high, capturing the sentiment with a comment, ”I didn’t expect that this very similar dataset was published even more recently than the dataset I found.” P22 21% participants discusse the utility of selecting ’Items per Page’, elaborating with a participant’s comment, ”I think displaying ten datasets at a time is very few, and I always have to click next page to refresh new datasets, which I don’t want to wait for.” N36 47% participants suggest the possible improvement of using colour tags for keyword matching, as multiple participants said, ”I would prefer these different colour tags to be about keyword information.” N37 37% participants signify the importance of ’Update Time’ and other dataset metadata like size and missing values ratio, summarised by a participant’s comment, ”as these factors influence whether the dataset warrants further examination.” N38 21% participants highlight the emphasis users put on ’Publisher’ and ’Updated’ in the expansion box, but found ’Contact,’ ’Identifiers,’ etc., less significant.
-One participant noted, ”I don’t need to understand the contact of a dataset at all, my primary purpose is to check whether the metadata of the dataset is suitable.” N39 42% mentions the confusion over non-English keywords and the difficulty in identifying query-related keywords from the list, summarised by one comment, ”This dataset covers too many keywords, and I just want to see at a glance which keywords from my search strings are included.” N40 16% participants expose user difficulty in understanding the ’Less Similar’ tag, with one participant stating, ”Instead of recording similarity here, why not directly state which keywords are included.” N41 16% participants found that datasets with the ”less similarity” tag were generally far from their target dataset.
-They found that only datasets with the ”similar” tag might be worth further exploration.
-4.2.4. Usage of Filters and Sorting
-P24 74% participants stated that the dual provision of filtering and sorting methods on the website was lauded for enhancing efficiency in dataset discovery.
-However, N51 58% participants exposed a time-consuming issue, where participants found that the page refreshes after each filter selection, especially problematic when applying multiple filters.
-N52 37% participants thought that refreshing the entire page after each filter selection could mislead users into thinking that the filter options are for single selection only, even though some filter options allow multiple selections, potentially resulting in imprecise search results.
-One participant, after selecting an option and noticing the page refresh and all expanded options collapse, skipped this filter and stated, ”It would be better if it was multi-select,” even though the filter was indeed multi-select.
-N53 16% participants spotlighted another issue where three participants were unaware that a filter was for single selection until they noticed that their previous selection was overwritten by the new one; a participant stated, ”I thought it was multi-select, and I wouldn’t have known it was single-select if the website’s upper right side didn’t display all my selected filter options.” N54 11% participants noted that the presence of numbers in the text of some filter options, such as ”Version 2”, can occasionally impact their interpretation of the number of datasets.
-One participant stated, ”At first glance, I thought there were only two datasets that met this filter requirement,” due to the confusion between the dataset count and the numerical information within the filter option’s text.
-Lastly, N55 26% participants found it easy to overlook the search box within each filter.
-One participant stated, ”as the placeholder text in the search box says ’Filter’, leading me to assume that it’s a description of the drop-down menu.”
-4.2.5. Dataset Actions
-P26 42% participants stated that the citation function was clearly visible and comprehensive, offering four different citation styles, which saved them significant time in creating citations.
-One participant stated, ”This has saved me a lot of time in creating citations.” N60 11% participants highlighted the absence of a direct LaTeX citation option, stating that manual LaTeX citation creation still constitutes a burdensome process.
-N61 58% participants stated that when they want to compare similarities between datasets, they need to manually copy and paste dataset information into a separate document, which they find cumbersome.
-One participant stated, ”It would be great if datasets could be compared directly on the website.”
-4.2.6. Feedback and Help
-P29: 68% participants indicated that they could easily locate the help documentation, and the information related to the functionality was easily discovered and understood.
-N63: 47% participants observed that while they can easily find general information like ”Contact Us” and ”Documentation”, they are often not interested in other options such as ”Publications” and ”Events” present in the navigation bar when searching for datasets.
-Their focus is primarily on dataset-related features.
-4.3. New Prototypes for Dataset Search Platform
-4.3.1. Initial Interaction
+4.1 Google Dataset Search
+In the initial interaction, 68% of participants found the search box, button, help, and feedback features accessible and user-friendly, describing the design as "neat and comfortable" (P1).
+79% appreciated the layout of the search results page, understanding the filters and praising their placement and the balance between preview and detailed information sections (P2).
+However, 26% noted the language switch button was hard to find at the bottom left, and 16% found its design confusing, suggesting the use of country flags or explicit labels (N1, N2).
+In multitasking scenarios, 21% found the detailed dataset information compressed and unreadable (N3).
+During the search process, 63% of participants used system suggestions to refine keywords, improving result relevance (P7).
+Google Dataset Search effectively corrected misspellings but sometimes failed to retrieve datasets for certain terms, offering spelling checks instead (P8, P9).
+21% effectively used Boolean and Exact Phrase operators, although specific prefixes did not yield desired results (P10, N14, N15).
+37% abandoned datasets due to missing values (N16), and the system sometimes misinterpreted queries (N17).
+Different spellings and sentence-like queries affected result relevance, and the Exclude Words operator did not function as expected (N18-N21).
+Searching for help with operators was unproductive as the User Guide page was missing, causing confusion (N22).
+In exploring datasets, participants focused on the title, update time, and dataset type, with a preference for content over the publisher (P13-P16).
+Ranking, especially in the top three listings, influenced choices.
+Participants valued descriptions, citation counts, and providers in the right-hand display section (P17).
+While 74% were indifferent to license specifics, 68% preferred its existence (P18).
+However, 53% initially cared about ranking, though 26% disregarded it and checked many datasets, sometimes scrolling to the page’s bottom (N29).
+Incomplete titles, limited preview information, and dense descriptions were common frustrations (N30-N32).
+37% highlighted a lack of detailed file information (N33), and despite difficulties in packing more information into previews, 16% still preferred a clean preview section (N34).
+After finding a suitable dataset, 21% clicked on ’explore at’ for more details, but 11% found this process cumbersome due to switching between multiple tabs (N35).
+Regarding filters and sorting, 68% intuitively understood filter names and effects (P23).
+However, 32% found it difficult to sort datasets by recency (N42).
+Multiple-choice filters like "topic" confused 26% of participants due to the misleading "+1" display, with 16% misinterpreting it as indicating only one topic selected (N43).
+16% preferred interactive animations for dropdown symbols, and 16% found the broad time spans in the ’Last updated’ filter problematic (N44, N45).
+47% were inconvenienced by filter resets upon search query modification (N46).
+Issues with datasets lacking usage permission tags and broad categorization within the ’Topics’ filter were noted (N47, N48).
+47% expressed confusion about dataset classification after applying filters (N49), and 11% ignored filters altogether due to their initial invisibility (N50).
+In terms of dataset actions, all participants easily found and used the "Save," "Share," and "Cite" buttons (P25).
+However, 16% preferred using their browser’s built-in share function (N56).
+When asked about sharing methods, 37% wouldn’t use Facebook and Twitter options, preferring work-related contexts (N57).
+Additionally, 21% found the dedicated email sharing feature less useful (N57).
+21% misunderstood the "Click to copy link" as informational text rather than an interactive button (N58), and 16% noted that the website does not specify the current citation format and standard (N59).
+Lastly, in feedback and help, 32% appreciated explanations on dataset search and provision (P27).
+84% easily found the "Help" and "Feedback" buttons, with some browsing the "Help" page content independently (P28).
+However, 11% encountered a ’404 Page not found’ error on the user guide page, increasing their reading time due to navigating multiple help pages (N62).
+4.2 data.europa.eu
+In the initial interaction with data.europa.eu, participants had mixed experiences.
+11% appreciated the comprehensive dataset summary and inclusion of datasets from 36 countries (P3).
+The page path display and window hover animations were positively noted by 26% and 42% of participants, respectively (P4, P5).
+However, 32% of users on smaller laptop screens found the search box hard to spot (N7), and 53% were distracted by unrelated windows and advertisements (N8).
+A UX designer highlighted the absence of search bar indicators (N9), and 37% found filter selection cumbersome due to shifting positions (N10).
+Some participants (11%) were frustrated by the inability to copy text from the preview section (N11), and color differentiation for tags was problematic for those with color blindness (N12).
+The ’show more’ button was seen as redundant by 11% of users (N13).
+During the search process, 63% found sorting by "Last modified" and "Last created" useful, especially under time constraints (P11).
+Separate search options for catalogues and datasets improved search efficiency for 21% of participants (P12).
+However, 26% noted that more search strings broadened the search scope instead of narrowing it (N23).
+Sorting by a single criterion led to extreme results for 16% (N24), and the absence of transition animations for dropdown symbols was noted by the same percentage of users (N25).
+In dataset exploration, 63% of participants appreciated the use of color tags for dataset types (P20), and 42% found the similar dataset feature beneficial (P21).
+The option to select ’Items per Page’ was useful for 21% (P22), although 47% suggested using color tags for keyword matching instead (N36).
+Metadata like ’Update Time’ and size were emphasized by 37% as important for dataset assessment (N37).
+Confusion over non-English keywords and the ’Less Similar’ tag was mentioned by 42% and 16% of participants, respectively (N39, N40).
+Some users found that datasets with the "less similarity" tag were not relevant (N41).
+Regarding filters and sorting, 74% praised the dual provision of filtering and sorting methods (P24).
+However, 58% found page refreshes after each filter selection time-consuming (N51).
+This led to confusion for 37%, as users thought filters were single-selection only (N52).
+The need for clearer indication of single-selection filters was noted by 16% of participants (N53), while 11% were confused by numbers in filter options’ text (N54).
+Additionally, 26% found the search box within each filter easy to overlook (N55).
+For dataset actions, 42% of participants found the citation function visible and comprehensive, offering multiple citation styles (P26).
+However, 11% missed a direct LaTeX citation option (N60), and 58% found comparing datasets cumbersome without a direct comparison feature on the website (N61).
+Lastly, in feedback and help, 68% of participants found the help documentation easy to locate and understand (P29).
+However, 47% noted that while general information like "Contact Us" and "Documentation" was easy to find, options such as "Publications" and "Events" were often irrelevant when focused on searching for datasets (N63).
+4.3 New Prototypes for Dataset Search Platform
+4.3.1 Initial Interaction
 We suggest that first impressions are crucial for dataset search platforms.
@@
 These design elements should be consistently implemented across various screen sizes to ensure a user-friendly experience.
-Figure 3. First Page of A Dataset Search Tool
-Prototype of First Page of Dataset Search Tool
-Figure 3. First Page of A Dataset Search Tool
-4.3.2. Search Process
+Figure 3: First Page of A Dataset Search Tool
+4.3.2 Search Process
 To enhance user interaction and platform efficiency, several improvements are recommended.
-Immediate auto-suggestions and spelling corrections could be provided through real-time prompts like ”did you mean.” Dataset search platform could also incorporate Natural Language Processing techniques such as entity extraction and synonym recognition for more accurate results.
+Immediate auto-suggestions and spelling corrections could be provided through real-time prompts like "did you mean." Dataset search platform could also incorporate Natural Language Processing techniques such as entity extraction and synonym recognition for more accurate results.
 Special attention should be paid to varying user needs; for instance, new users might appreciate a help query option in the search bar, while other users may prioritize datasets based on factors like geographic location (Shown in Figure  4).
-Additionally, implementing common search operators like ”AND,” ”OR,” and ”NOT” can refine search specificity.
+Additionally, implementing common search operators like "AND," "OR," and "NOT" can refine search specificity.
 To make these operators more accessible, they could be visualized in the search bar interface, accompanied by dropdown menus or clickable icons.
@@
 This approach not only informs users about advanced search tools but also enables them to generate highly targeted results from the beginning (Shown in Figure  5).
-Figure 4. Search Bar Considering Varied User Needs
-Prototype of Search Bar Considering Varied User Needs
-Figure 4. Search Bar Considering Varied User Needs
-Figure 5. Search Bar Considering Enhanced Precision
-Prototype of Search Bar Considering Enhanced Precision
-Figure 5. Search Bar Considering Enhanced Precision
-4.3.3. Dataset Exploration
+Figure 4: Search Bar Considering Varied User Needs
+Figure 5: Search Bar Considering Enhanced Precision
+4.3.3 Dataset Exploration
 Dataset previews need to be brief but informative.
@@
 The previews should also be succinct while offering depth, including expandable segments for extra details.
-These segments can feature key terms, descriptions of columns, and ratios of missing values, presented in a bullet-point format (See Figure  7).
-Figure 6. User-Focused Dataset Previews Section
-Prototype of Dataset Preview Considering Prioritizing User Preferences
-Figure 6. User-Focused Dataset Previews Section
-Figure 7. Expandable Section For Additional Information
-Expandable Section For Additional Information about dataset metadata
-Figure 7. Expandable Section For Additional Information
-Figure 8. Simplified Result Display Section
-Prototype of Optimizing Result Previews for Enhanced User Experience
-Figure 8. Simplified Result Display Section
+These segments can feature key terms, descriptions of columns, and ratios of missing values, presented in a bullet-point format.
+Figure 6: User-Focused Dataset Previews Section
+Figure 7: Simplified Result Display Section
 Given the numerous dataset previews, users can face reading fatigue.
 To counteract this, a colour-coded scheme could be employed to highlight datasets meeting user-defined criteria.
-For example, using dark green to signify criteria that match the user’s definitions, intermediate blue for criteria that don’t match, and pale orang for criteria that the user hasn’t defined (Shown in Figure  8), while being accessible to colour-blind users, tested in Section  4.3.5.
-4.3.4. Usage of Filters and Sorting
+For example, using dark green to signify criteria that match the user’s definitions, intermediate blue for criteria that don’t match, and pale orang for criteria that the user hasn’t defined (Shown in Figure  7), while being accessible to colour-blind users, tested in Section  4.3.5.
+4.3.4 Usage of Filters and Sorting
 To enhance the user experience across platforms, several improvements are suggested.
@@
 Ideally positioned at the top of the page, filters should not necessitate a full page refresh upon each selection.
-Instead, a subtle bottom-right corner notification should indicate the number of datasets that fit the selected criteria (Shown in Figure  9).
-Figure 9. Improved Filter Design
-Prototype of Filters Considering Optimizing Placement and Interactivity
-Figure 9. Improved Filter Design
+Instead, a subtle bottom-right corner notification should indicate the number of datasets that fit the selected criteria (Shown in Figure  8).
+Figure 8: Improved Filter Design
 Sorting should employ dual criteria: initial sorting based on relevance, followed by secondary sorting according to publication date.
-These improvements aim for a more efficient and user-friendly dataset search process, with subsequent figures outlining the proposed changes (Shown in Figure  10).
-Figure 10. Dual Criteria Sorting
-Prototype of Dual Criteria Sorting
-Figure 10. Dual Criteria Sorting
-4.3.5. Dataset Actions
+These improvements aim for a more efficient and user-friendly dataset search process.
+4.3.5 Dataset Actions
 Improvements in dataset actions could be multifaceted.
@@
 A dark green fill could signify attributes where another dataset performs better than the currently selected one, a intermediate blue fill for areas where the current dataset excels, and a pale orange fill to highlight unique features not offered by the selected dataset.
-This visual guide would streamline the decision-making process, rendering it more intuitive and less cognitively taxing for users (Shown in Figure  11).
-Figure 11. Datasets Comparison Page
-Prototype of Datasets Comparison Page
-Figure 11. Datasets Comparison Page
-The researcher used the Coblis Colour Blindness Simulator (col, 2023) to test 8 types of colour blindness on the Figure  11 provided .
-The results show clear variations in colour fills for each type of colour blindness, ensuring accessibility for those with colour vision deficiencies (Shown in Figure  12).
-Figure 12. Eight Colour Blindness Tests
-Coblis Colour Blindness Simulator
-Figure 12. Eight Colour Blindness Tests
-4.3.6. Feedback and Help
+This visual guide would streamline the decision-making process, rendering it more intuitive and less cognitively taxing for users (Shown in Figure  9).
+Figure 9: Datasets Comparison Page
+The researcher used the Coblis Colour Blindness Simulator [1] to test 8 types of colour blindness on the Figure  9 provided .
+The results show clear variations in colour fills for each type of colour blindness, ensuring accessibility for those with colour vision deficiencies (Shown in Figure  10).
+Figure 10: Eight Colour Blindness Tests
+4.3.6 Feedback and Help
 For enhancements, both platforms could incorporate introductory tutorials to familiarize new users with their respective features.
 A step-by-step guide, necessitating users to acknowledge each tutorial segment before advancing, can ensure a more comprehensive understanding of the platform’s functionalities.
-By instituting this level of guidance, platforms can increase user engagement, reduce initial learning curves, and subsequently, elevate the overall user experience (Shown in Figure  13).
-Figure 13. Tutorial Illustration
-Prototype of Tutorial Illustration
-Figure 13. Tutorial Illustration
-5. Discussion and Limitations
+By instituting this level of guidance, platforms can increase user engagement, reduce initial learning curves, and subsequently, elevate the overall user experience (Shown in Figure  11).
+Figure 11: Tutorial Illustration
+5 Discussion and Limitations
 Our study found that users experience significant reading fatigue due to the way metadata is presented on dataset search websites.
@@
 Our primary focus was on exploring the accessibility of dataset metadata, specifically how to present it to improve search efficiency, precision, and reduce reading fatigue, addressing key user experience challenges today.
-6. Conclusion and Future Works
+6 Conclusion and Future Works
 This research assesses user interactions in dataset search on two platforms: Google Dataset Search and data.europa.eu.
@@
 Physical eye-tracking devices could offer a solution to the limitations encountered with software-based approaches, as they are less likely to be hindered by website security protocols.
-Acknowledgements.
-The author wishes to acknowledge the contributions of supervisors Professor Elena Simperl and Dr.
-Albert Meroño Peñuela for their insightful guidance and inspiration.
-Thanks are also directed to Bohui Zhang for editorial assistance.
-Gratitude is extended to the varied participants of the user study and the Department of Informatics at King’s College London for providing essential research materials and support in participant recruitment.